OpenStudy (anonymous):

Draw a line of best fit on your scatter plot. Choose the best prediction for how many calories from fat you would expect to find in a food that contains 25 grams of fat?

OpenStudy (amistre64):

\[f(x)=\frac{n\sum(xy)-\sum x\sum y}{n\sum(xx)-\sum x\sum x}(X-\bar X)+\bar Y\]